# A Multimodal Flow Foundation Model for 3D Molecules and Materials

### v1

@@ Abstract
General-purpose 3D chemical modeling encompasses molecules and materials, requiring both generative and predictive capabilities. However, most existing AI approaches are optimized for a single domain (molecules or materials) and a single task (generation or prediction), which limits representation sharing and transfer. We introduce Zatom-1, the first foundation model that unifies generative and predictive learning of 3D molecules and materials. Zatom-1 is a Transformer trained with a multimodal flow matching objective that jointly models discrete atom types and continuous 3D geometries. This approach supports scalable pretraining with predictable gains as model capacity increases, while enabling fast and stable sampling. We use joint generative pretraining as a universal initialization for downstream multi-task prediction of properties, energies, and forces. Empirically, Zatom-1 matches or outperforms specialized baselines on both generative and predictive benchmarks, while reducing the generative inference time by more than an order of magnitude. Our experiments demonstrate positive predictive transfer between chemical domains from joint generative pretraining: modeling materials during pretraining improves molecular property prediction accuracy.
@@ 1 Introduction
Recently, deep learning has begun to unify the generation of new 3D molecules and materials (Joshi et al., 2025), and several specialized prediction models currently exist for advanced molecule and material property prediction (Liao and Smidt, 2023; Qu and Krishnapriyan, 2024). Nonetheless, no prior works offer a unified model capable of both generative modeling and representation learning of 3D molecules and materials. To build such a unified model, several design choices must be addressed, including how to tokenize its inputs as well as what and how to learn using these tokens.
@@ 2.4 Stage 2: Finetuning
As illustrated in Figure 1, we treat conditional flow matching of the modalities $\left(\bm{A},\bm{X},\bm{F},\bm{L}_{\text{len}},\bm{L}_{\text{ang}}\right)$ as a self-supervised pretraining task akin to bidirectional modeling used to pretrain non-autoregressive BERT-style models (Devlin et al., 2019) for enhanced representation learning (Zhang et al., 2025) and label-free transfer learning. Once the model weights listed up to Line 10 of Algorithm 1 have been (pre)trained, for finetuning, we freeze these weights and enable training of the $M$-layer auxiliary Transformer trunk weights following Line 10 using trunk layer $K$’s frozen embeddings. By default, $K=L$, which we ablate in Section 4.
@@ 5 Conclusion
In this work, we introduced Zatom-1, the first general-purpose foundation model for 3D molecules and materials. Empirical results show that Zatom-1 achieves state-of-the-art molecule generation and competitive materials generation, with fast training and inference speeds. Additionally, it demonstrates the advantages of generative pretraining for downstream prediction tasks, marking SciML’s first instance of positive predictive transfer learning through joint generative pretraining. As a single neural network capable of multiple fundamental tasks in molecular and material science, Zatom-1’s code and weights are freely available.
@@ Appendix A Related Work
Overall, prior work in molecular/material generation has focused either on discrete (graph/SMILES) models or on separate generative flows/diffusions for molecules or crystals (Hoogeboom et al., 2022; Xie et al., 2022). Our approach differs by unifying generative and predictive modeling for all-atom 3D chemistry. We draw on advances in flow matching and multimodal architecture design (Song et al., 2023; Miller et al., 2024; Vonessen et al., 2025), and we treat generative modeling itself as a pretraining strategy. This differs from previous scientific foundation models that often focus on language or graph representations (Pyzer-Knapp et al., 2025; Choi et al., 2025; Wadell et al., 2025). By explicitly modeling the ambient 3D space of atoms and elements together, we extend the flow-based generative paradigm to create a single foundation model across 3D molecular and material domains.
@@ B.2 Empirical results
Test (% pass) $\uparrow$ Symphony (QM9-only) Eq. Diff. (QM9-only) ADiT (jointly trained) Platom-1 (QM9-only) Zatom-1 (QM9-only) Zatom-1 (jontly trained) Atoms connected 99.92 99.88 99.70 99.97 99.95 99.98 Bond angles 99.56 99.98 99.85 99.94 99.90 99.95 Bond lengths 98.72 100.00 99.41 99.96 99.89 99.97 Aromatic ring flat 100.00 100.00 100.00 100.00 100.00 100.00 Double bond flat 99.07 98.58 99.98 100.00 100.00 99.99 Internal energy 95.65 94.88 95.86 99.73 99.63 99.78 No steric clash 98.16 99.79 99.79 99.84 99.81 99.81

### v2

@@ Abstract
General-purpose 3D chemical modeling encompasses molecules and materials, requiring both generative and predictive capabilities. However, most existing AI approaches are optimized for a single domain (molecules or materials) and a single task (generation or prediction), which limits representation sharing and transfer. We introduce Zatom-1, the first end-to-end, fully open-source foundation model that unifies generative and predictive learning of 3D molecules and materials. Zatom-1 is a Transformer trained with a multimodal flow matching objective that jointly models discrete atom types and continuous 3D geometries. This approach supports scalable pretraining with predictable gains as model capacity increases, while enabling fast and stable sampling. We use joint generative pretraining as a universal initialization for downstream multi-task prediction of properties, energies, and forces. Empirically, Zatom-1 matches or outperforms specialized baselines on both generative and predictive benchmarks, while reducing the generative inference time by more than an order of magnitude. Our experiments demonstrate positive predictive transfer between chemical domains from joint generative pretraining: modeling materials during pretraining improves molecular property prediction accuracy.
@@ 1 Introduction
Recently, deep learning has begun to unify the generation of new 3D molecules and materials (Joshi et al., 2025; Zhang et al., 2025a), and several specialized prediction models currently exist for advanced molecule and material property prediction (Liao and Smidt, 2023; Qu and Krishnapriyan, 2024). Nonetheless, no prior works offer an end-to-end, fully open-source model capable of both generative modeling and representation learning of 3D molecules and materials. To build such a unified model, several design choices must be addressed, including how to tokenize its inputs as well as what and how to learn using these tokens.
@@ 2.4 Stage 2: Finetuning
As illustrated in Figure 1, we treat conditional flow matching of the modalities $\left(\bm{A},\bm{X},\bm{F},\bm{L}_{\text{len}},\bm{L}_{\text{ang}}\right)$ as a self-supervised pretraining task akin to bidirectional modeling used to pretrain non-autoregressive BERT-style models (Devlin et al., 2019) for enhanced representation learning (Zhang et al., 2025b) and label-free transfer learning. Once the model weights listed up to Line 10 of Algorithm 1 have been (pre)trained, for finetuning, we freeze these weights and enable training of the $M$-layer auxiliary Transformer trunk weights following Line 10 using trunk layer $K$’s frozen embeddings. By default, $K=L$, which we ablate in Section 4.
@@ 5 Conclusion
In this work, we introduced Zatom-1, the first general-purpose foundation model for 3D molecules and materials. Empirical results show that Zatom-1 achieves state-of-the-art molecule generation and competitive materials generation, with fast training and inference speeds. Additionally, it demonstrates the advantages of generative pretraining for downstream prediction tasks, marking SciML’s first instance of positive predictive transfer learning through joint generative pretraining. As a single neural network capable of multiple fundamental tasks in molecular and material science, to advance the development of atomic foundation models, Zatom-1’s code and weights are freely available at https://github.com/Zatom-AI/zatom.
@@ Appendix A Related Work
Overall, prior work in molecular/material generation has focused either on discrete (graph/SMILES) models or on separate generative flows/diffusions for molecules or crystals (Hoogeboom et al., 2022; Xie et al., 2022). Our approach differs by unifying generative and predictive modeling for all-atom 3D chemistry. We draw on advances in flow matching and multimodal architecture design (Song et al., 2023; Miller et al., 2024; Vonessen et al., 2025), and we treat generative modeling itself as a pretraining strategy. This differs from previous scientific foundation models that often focus on language or graph representations (Pyzer-Knapp et al., 2025; Choi et al., 2025; Wadell et al., 2025), or that rely on autoregression and do not provide end-to-end experimental reproducibility (Zhang et al., 2025a). By explicitly modeling the ambient 3D space of atoms and elements together, we extend the flow-based generative paradigm to create a single foundation model across 3D molecular and material domains.
@@ B.2 Empirical results
Test (% pass) $\uparrow$ Symphony (QM9-only) Eq. Diff. (QM9-only) ADiT (jointly trained) Platom-1 (QM9-only) Zatom-1 (QM9-only) Zatom-1 (jointly trained) Atoms connected 99.92 99.88 99.70 99.97 99.95 99.98 Bond angles 99.56 99.98 99.85 99.94 99.90 99.95 Bond lengths 98.72 100.00 99.41 99.96 99.89 99.97 Aromatic ring flat 100.00 100.00 100.00 100.00 100.00 100.00 Double bond flat 99.07 98.58 99.98 100.00 100.00 99.99 Internal energy 95.65 94.88 95.86 99.73 99.63 99.78 No steric clash 98.16 99.79 99.79 99.84 99.81 99.81